# Self-Supervised Learning and Opportunistic Inference for Continuous Monitoring of Freezing of Gait in Parkinson’s Disease

### v1

@@ 3.2 Opportunistic Inference
Figure 2: Model Activation Module: Activity threshold-based triggering mechanism for computational offloading and battery life extension in wearable FoG detection system.
@@
$\left\{\begin{array}[]{ll}1\text{ (active)}&\text{if }M_{i}\geq\alpha\\ 0\text{ (inactive)}&\text{otherwise}\end{array}\right.$ (1)
@@
3.3 Imbalanced Training Data Mitigation
@@
(a) DHWT: Overlapping size increases.
(b) Balanced class distribution using DHWT
Figure 3: DHWT segmentation process for training set (a); FoG proportion using standard vs. DHWT segmentation (b).
For instance, in our experimental evaluation shown in Fig. 3a, we applied a 50% overlap for non-FoG periods and a 75% overlap for FoG episodes during training. During inference, windows are segmented using standard, non-variable overlaps to simulate real-world conditions where FoG episodes are not pre-identified. Using fixed-length overlaps typically results in an unbalanced training set, as illustrated in the left bar of Fig. 3b, with 63.3% non-FOG and 36.7% FOG instances. In contrast, the DHWT segmentation approach achieves a more balanced distribution, as shown in the right bar of the same figure, with non-FOG and FOG data at 45% and 55%, respectively. For test set generation, standard segmentation with a fixed 50% overlap (advancing every 1.5 seconds) mimics actual operating conditions, processing data from the preceding 3 seconds.
3.4 Dataset
The dataset for this study is the publicly available tDCS FoG dataset, comprising movement data from PD subjects in both medicated (‘On’) and unmedicated (‘Off’) states during Freezing of Gait (FoG) provocation protocols [21]. Data were collected using a 3D accelerometer attached to the lower back, recording at 128 Hz, resulting in 1132 FoG episodes (285 minutes total) and 15.3 hours of recording. Each FoG episode was videotaped and annotated by experts. The demographic and clinical characteristics of the subjects are summarized in Table 2. The dataset contains events labeled as “Normal” or “Freezing of Gait” (FoG). The distribution of these events is summarized in Table 3. The sessions followed the FoG provocation protocol as described in seminal studies by Reches et al. [22] and Manor et al. [23], and originally defined by Ziegler et al. [24].
Table 2: Summary of patient characteristics
@@ 3.4 Dataset
Years Since DX mean(SD), [min, max] 10.5 (5.9), [1, 23]
Table 3: Event distribution during FoG provocation trials
@@
Normal 28.01 35.41 63.42
FoG 18.73 17.86 36.59
We resampled the data to 40 Hz, which is considered an effective frequency for recognizing human activity through accelerometers in both healthy individuals and PD patients, including those experiencing Freezing of Gait (FoG) episodes. This frequency captures the typical freeze band (3-8 Hz) while reducing memory load and computational complexity.
Input: $X$: Set of inference data frames,
$[\alpha_{start},\alpha_{final}]$: Initial and final threshold,
$P_{0}$: Baseline performance metric,
$f(.)$: Inference Model,
$A(.)$: returns magnitude for $x_{i}\in X$
Output: $\alpha_{opt}:$ Optimal activity threshold,
$X_{opt}$: Active frames for ($\alpha_{opt}$)
Initialization $\alpha\leftarrow\alpha_{start}$
while $\alpha\leq\alpha_{final}$ do
       $X_{active}=\{x_{i}\in X\mid A(x_{i})\geq\alpha\}$triangleright Set of active frames;
      $N_{\alpha}\leftarrow{}\mid X_{active}\mid$triangleright Size of active frames;
      $P\leftarrow f(X_{\alpha})$
      if  $\mid P-P_{0}\mid\leq\theta$  then
             $\alpha\leftarrow\alpha+\delta\alpha$
      else
             $\alpha_{opt}\leftarrow\alpha$;
             $X_{opt}=\{x_{i}\in X\mid A(x_{i})\geq\alpha_{opt}\}$ ;
             break;
       end if
end while
return $\textbf{return }\alpha_{opt},X_{opt}$
Algorithm 1 Activity Threshold Optimization Algorithm for Prolonged Battery Life
3.5 Mean Value Removal and Labeling
@@
3.6 Leave One Group Out (LOGO)
@@
3.7 Model Development & Experimental Setup
@@
Figure 4: Architecture of stacked 1D CNN model. Input to the model is 3-minute raw sensor data.
@@
3.8 Performance Measures
@@
Figure 5: A: FoG Episode, B: False Positive, C: Normal Episode, D: Detected FoG Episode (TP), E: Detected Normal Episode (TN), L: Latency
Sensitivity/Recall measures the ratio of correctly detected FOG windows and specificity measures the ratio of correctly identified non-FoG windows. Precision evaluates the model’s ability to avoid false positives. The F1 score is the harmonic mean of sensitivity and precision which is used to assess performance on the imbalanced dataset [25].
3.9 FoG Detection Post-processing
@@
Table 4: Performance of proposed LIFT-PD framework. Key Metrics Defined: PREC: Precision, REC: Recall, SENS: Sensitivity, ACC: Accuracy, SPEC: Specificity, DFE: Detected FoG Episodes, DFW: Detected FoG Windows, STD: Standard Deviation. Performance metrics for the baseline supervised model are presented within parentheses ().
@@
3.9.1 FoG Episode
In our study, we assessed the detection of FoG episodes—defined as consecutive windows labeled as FoG—by computing both the percentage of episodes detected and the proportion of FoG accurately identified within each episode. An episode is considered detected if at least one window within it is correctly identified. FoG episodes were categorized into three groups: short ($<6$ seconds), medium ($6-12$ seconds), and long ($>12$ seconds).
@@ 3.9.2 Computational Complexity
Table 5: Performance comparison across diverse groups. Performance metrics for the baseline supervised model are presented within parentheses ().
@@
Gender (40) Female 0.78 (0.67) 0.75 (0.77) 0.76 (0.72) 0.84 (0.71) 0.80 (0.77)
Male (32) 0.71 (0.74) 0.67 (0.64) 0.69 (0.68) 0.83 (0.86) 0.77 (0.77)
Age (40) Old (20) 0.80 (0.75) 0.71 (0.71) 0.75 (0.74) 0.86 (0.82) 0.80 (0.78)
Mid-age 0.70 (0.66) 0.80 (0.79) 0.74 (0.72) 0.82 (0.79) 0.81 (0.79)
Medication On 0.82 (0.78) 0.79 (0.72) 0.81 (0.75) 0.84 (0.82) 0.86 (0.81)
Off 0.78 (0.70) 0.8 (0.74) 0.79 (0.72) 0.79 (0.72) 0.82 (0.78)
Random (40) 1 (20) 0.66 (0.61) 0.82 (0.82) 0.73 (0.70) 0.81 (0.75) 0.81 (0.77)
2 (20) 0.82 (0.77) 0.86 (0.86) 0.84 (0.81) 0.83 (0.78) 0.84 (0.82)
4 Results
4.1 Performance Analysis
@@
Figure 6: Receiver Operating Characteristic (ROC) curves for supervised and self-supervised models
@@
4.1.1 Generalizibility of DHWT
Table 6: Performance comparison of DHWT across datasets. The second row shows relative improvement over the baseline.
@@ 4.1.1 Generalizibility of DHWT
tdcs 1.8s $\uparrow$880.94s (0.92s) 0.74 $\uparrow$7.20.8 $\uparrow$5.30.82 $\uparrow$6.50.83 $\uparrow$5.1%
Daphnet 0.14s $\uparrow$420.134s (0.132s) 0.55 $\uparrow$1610.44 $\uparrow$420.94 $\uparrow$1080.87 $\uparrow$71%
MotionSense 0.08 $\uparrow$1420.022s (0.024s) 1 $\uparrow$3.10.99 $\uparrow$2.51 $\uparrow$1.20.99 $\uparrow$1.22%
We expanded our analysis to Daphnet [26] and two activities from MotionSense [27] (as a binary classification), in addition to the primary tdcs dataset. The results show an increase in training time across the datasets due to class imbalance handling, while the inference time remained stable. DHWT significantly improved precision, F1-score, and specificity in datasets with class imbalance, confirming its effectiveness across diverse conditions. Daphnet showed a 161% boost in precision and a 108% improvement in specificity. These results support the generalizability of DHWT without compromising computational efficiency (Table 6).
4.1.2 Robustness Across Diverse Subjects
@@
4.2 Compare with State-of-the-Art Models
Table 7 compares the performance of our LIFT-PD model with state-of-the-art methods for detecting FoG in PD patients. Although the Multi-head CNN [12] achieves the highest detection rates for FoG episodes (97.27%) and windows (94.64%), its precision (0.545) and specificity (0.491) are low, indicating a high false positive rate. High false positives can reduce the effectiveness of cueing due to patients’ adaptiveness to “always on” interventions [28] [29]. The one-class classifier [10] shows high precision (0.856) and specificity (0.891) but lower recall (0.716), missing many true FoG events, which can lead to inadequate monitoring and delayed interventions. The semi-supervised model [11] shows a recall around 12% lower and specificity approximately 13% lower than LIFT-PD, making it less suitable for in-home monitoring with a single accelerometer.
Table 7: Comparison with state-of-the-art models

### v2

@@ 3.2. Opportunistic Inference
Figure 2. Model Activation Module: Activity threshold-based triggering mechanism for computational offloading and battery life extension in wearable FoG detection system.
@@
(3) $\left\{\begin{array}[]{ll}1\text{ (active)}&\text{if }M_{i}\geq\alpha\\ 0\text{ (inactive)}&\text{otherwise}\end{array}\right$.
@@
3.3. Imbalanced Training Data Mitigation
@@
Figure 3. DHWT segmentation process for training set (a); FoG proportion using standard vs. DHWT segmentation (b).
For instance, in our experimental evaluation shown in Fig. LABEL:subfig:dhwt, we applied a 50% overlap for non-FoG periods and a 75% overlap for FoG episodes during training. During inference, windows are segmented using standard, non-variable overlaps to simulate real-world conditions where FoG episodes are not pre-identified. Using fixed-length overlaps typically results in an unbalanced training set, as illustrated in the left bar of Fig. LABEL:subfig:dhw-samples, with 63.3% non-FoG and 36.7% FoG instances. In contrast, the DHWT segmentation approach achieves a more balanced distribution, as shown in the right bar of the same figure, with non-FoG and FoG data at 45% and 55%, respectively. For test set generation, standard segmentation with a fixed 50% overlap (advancing every 1.5 seconds) mimics actual operating conditions, processing data from the preceding 3 seconds. A detailed explanation of the choice of a 3-second window length is provided in Section 4.7. The selection of 75% overlap for FoG and 50% for non-FoG segments was guided by both empirical validation and domain expertise. With a 3-s window and typical gait-cycle durations of 1.2–1.5s in PD patients (yin2024gait; pistacchi2017gait), a 75% overlap (hop $\approx$ 0.75 s) ensures each training window contains multiple stride transitions and captures the temporal evolution of a FoG bout. In contrast, a 50% overlap (hop $\approx$ 1.5 s) during non-FoG activity reduces sample redundancy and computation. These ratios balance class representation ($\approx$ 1:1 FoG:non-FoG windows) while maintaining physiologically meaningful temporal coverage consistent with expert-observed gait dynamics.
3.4. Dataset
The dataset for this study is the publicly available tDCS FoG dataset, comprising movement data from PD subjects in both medicated (‘On’) and unmedicated (‘Off’) states during Freezing of Gait (FoG) provocation protocols (howard2023parkinsons). Data were collected using a 3D accelerometer attached to the lower back, recording at 128 Hz, resulting in 1132 FoG episodes (285 minutes total) and 15.3 hours of recording. Each FoG episode was videotaped and annotated by experts (howard2023parkinsons). The demographic and clinical characteristics of the subjects are summarized in Table 2. The dataset contains events labeled as “Normal” or “Freezing of Gait” (FoG). The distribution of these events is summarized in Table 3. The sessions followed the FoG provocation protocol as described in seminal studies by Reches et al. (reches2020) and Manor et al. (manor2021multitarget), and originally defined by Ziegler et al. (ziegler2010).
Table 2. Summary of patient characteristics
@@ 3.4. Dataset
Years Since DX
mean(SD), [min, max]
10.5 (5.9), [1, 23]
Table 3. Event distribution during FoG provocation trials
@@
Normal
28.01 35.41 63.42
FoG
18.73 17.86 36.59
Figure 4. Density distribution and box plot of the Unified Parkinson’s Disease Rating Scale (UPDRS) scores, separated by sex. The top panel displays the density distribution of UPDRS scores, with males shown in blue and females in purple. The bottom panel presents a box plot indicating variations in UPDRS scores between sexes, highlighting the median scores, interquartile ranges, and outliers.
We resampled the data to 40 Hz, which is considered an effective frequency for recognizing human activity through accelerometers in both healthy individuals and PD patients, including those experiencing Freezing of Gait (FoG) episodes. This frequency captures the typical freeze band (3-8 Hz) while reducing memory load and computational complexity. This reduced frequency is sufficient for capturing the typical freeze band (ranging from 3 to 8 Hz) that appears during FoG episodes. Additionally, downsampling to 40 Hz helps to decrease the memory load and computational complexity of the data processing.
The density plot (Fig 4, top panel) suggests that while both distributions are roughly normal, males exhibit a higher peak density in the score range of 30-40, compared to females who show a broader peak around 40-50. The box plot (Figure 4, bottom panel) indicates a slightly lower median UPDRS score for males compared to females, although overlaps in the interquartile ranges and the presence of outliers in both groups suggest variability within the scores.
3.5. Mean Value Removal and Labeling
@@
3.6. Leave One Group Out (LOGO)
@@
3.7. Model Development & Experimental Setup
@@
Figure 5. Architecture of stacked 1D CNN model. Input to the model is 3-minute raw sensor data.
@@
Algorithm 1 Activity Threshold Optimization (ATO) Algorithm for Prolonged Battery Life
1:Input: $X$: Set of inference data frames, $\alpha_{start},\alpha_{final}]$: Initial and final threshold, $P_{0}$: Baseline performance metric, $f(\cdot)$: Inference Model, $A(\cdot)$: returns magnitude for $x_{i}\in X$
2:Output: $\alpha_{opt}:$ Optimal activity threshold, $X_{opt}$: Active frames for ($\alpha_{opt}$)
3:Begin
4:initialize $\alpha\leftarrow\alpha_{start}$
5:while $\alpha\leq\alpha_{final}$ do
6:  $X_{active}=\{x_{i}\in X\mid A(x_{i})\geq\alpha\}$ $\triangleright$ Set of active frames
7:  $N_{\alpha}\leftarrow{}\mid X_{active}\mid$ $\triangleright$ Size of active frames
8:  $P\leftarrow f(X_{\alpha})$
9:  if $|P-P_{0}|\leq\theta$ then
10:   $\alpha\leftarrow\alpha+\delta\alpha$
11:  else
12:   $\alpha_{\text{opt}}\leftarrow\alpha$
13:   $X_{opt}\leftarrow\{x_{i}\in X\mid A(x_{i})\geq\alpha_{opt}\}$
14:   break
15:  end if
16:end while
17:return $\alpha_{opt},X_{opt}$ $\triangleright$ optimal threshold, activity frames
3.8. Performance Measures
@@
Figure 6. A: FoG Episode, B: False Positive, C: Normal Episode, D: Detected FoG Episode (TP), E: Detected Normal Episode (TN), L: Latency
Sensitivity/Recall measures the ratio of correctly detected FoG windows and specificity measures the ratio of correctly identified non-FoG windows. Precision evaluates the model’s ability to avoid false positives. The F1 score is the harmonic mean of sensitivity and precision which is used to assess performance on the imbalanced dataset (10054906).
3.9. FoG Detection Post-processing
@@
Table 4. Performance of proposed LIFT-PD framework. Key Metrics Defined: PREC: Precision, REC: Recall, SENS: Sensitivity, ACC: Accuracy, SPEC: Specificity, DFE: Detected FoG Episodes, DFW: Detected FoG Windows, STD: Standard Deviation. Performance metrics for the baseline supervised model are presented within parentheses ().
@@
3.9.1. FoG Episode
In our study, we assessed the detection of FoG episodes—defined as consecutive windows labeled as FoG—by computing both the percentage of episodes detected and the proportion of FoG accurately identified within each episode. An episode is considered detected if at least one window within it is correctly identified. FoG episodes were categorized into three groups: short (<6<6 seconds), medium (6126-12 seconds), and long (>12>12 seconds).
@@ 3.9.2. Computational Complexity
Table 5. Performance comparison across diverse groups. Performance metrics for the baseline supervised model are presented within parentheses ().
@@
Gender
Female
0.78 (0.67)
0.75 (0.77)
0.76 (0.72)
0.84 (0.71)
0.80 (0.77)
Male (32)
0.71 (0.74)
0.67 (0.64)
0.69 (0.68)
0.83 (0.86)
0.77 (0.77)
Age
Old (20)
0.80 (0.75)
0.71 (0.71)
0.75 (0.74)
0.86 (0.82)
0.80 (0.78)
Mid-age
0.70 (0.66)
0.80 (0.79)
0.74 (0.72)
0.82 (0.79)
0.81 (0.79)
Medication
On
0.82 (0.78)
0.79 (0.72)
0.81 (0.75)
0.84 (0.82)
0.86 (0.81)
Off
0.78 (0.70)
0.8 (0.74)
0.79 (0.72)
0.79 (0.72)
0.82 (0.78)
Random
1 (20)
0.66 (0.61)
0.82 (0.82)
0.73 (0.70)
0.81 (0.75)
0.81 (0.77)
2 (20)
0.82 (0.77)
0.86 (0.86)
0.84 (0.81)
0.83 (0.78)
0.84 (0.82)
4. Results
4.1. Performance Analysis
@@
Figure 7. Receiver Operating Characteristic (ROC) curves for supervised and self-supervised models
@@
Table 6. Comparison between the proposed Self-Supervised Learning (SSL) model and two Semi-Supervised Learning baselines—Mean-Teacher (MT) and $\Pi$ model—using the same 1-D CNN backbone. Metrics are averaged across groups; episode-level results (DFE) show detection rates for different FoG episode durations.
Group Precision Rec/Sens/DFW F1 Score Accuracy Specificity
SSL MT $\Pi$ SSL MT $\Pi$ SSL MT $\Pi$ SSL MT $\Pi$ SSL MT $\Pi$
1 0.66 0.65 0.66 0.82 0.75 0.73 0.73 0.70 0.70 0.81 0.79 0.80 0.81 0.81 0.84
2 0.82 0.76 0.80 0.86 0.74 0.81 0.84 0.75 0.80 0.84 0.77 0.82 0.83 0.80 0.82
Average 0.74 0.71 0.74 0.84 0.75 0.77 0.79 0.72 0.75 0.83 0.78 0.81 0.82 0.81 0.83
% of Detected FoG Episodes (DFE)
Small (0–6 s) Medium (6–12 s) Large (>>12 s)
SSL MT $\Pi$ SSL MT $\Pi$ SSL MT $\Pi$
83.3% 77.7% 79.7% 100% 96.0% 97.0% 100% 99.2% 99.2%
4.1.1. Compare with Baselines
We also compared the performance of our self-supervised learning (SSL) approach with two semi-supervised learning (Semi-SL) baselines for FoG detection. To isolate the benefit of our representation-learning strategy, we replaced the SSL pre-training stage in LIFT-PD with (i) a $\Pi$‑model–based semi-supervised pipeline (FixMatch) and (ii) a Mean-Teacher (MT) variant (tarvainen2017mean), while keeping the backbone network, DHWT segmentation, and opportunistic inference module identical across models. Both Semi-SL methods leverage unlabeled data through consistency regularization, yet differ in how pseudo-labels are formed: the $\Pi$-model relies on confidence-filtered student predictions, whereas the MT uses an exponential-moving-average (EMA) teacher to provide soft, stable targets.
Although both baselines improved over purely supervised training, they required a larger fraction of labeled data ($\approx$65–75% of the training set, vs. 40% for LIFT‑PD) for stable pseudo-label convergence. Due to the sparsity and variability of FoG episodes, early pseudo-labels in the $\Pi$-model tended to be noisy, while the MT showed better stability but lower sensitivity under limited label ratios. In contrast, our SSL pre-text task learns label-agnostic temporal structure by reconstructing masked sensor segments, enabling the encoder to generalize from all recordings without relying on tentative labels. As a result, LIFT-PD achieved higher recall (0.84 vs. 0.75 [MT] / 0.77 $\Pi$]) and F1-score (0.79 vs. 0.72 [MT] / 0.75 $\Pi$]), while matching or exceeding both baselines in precision, accuracy, and specificity (Table 6).
4.1.2. Generalizibility of DHWT
Table 7. Performance comparison of DHWT across datasets. The second row shows relative improvement over the baseline.
@@ 4.1.2. Generalizibility of DHWT
tdcs
1.8s
$\uparrow$88%
0.94s
0.74
$\uparrow$7.2%
0.8
$\uparrow$5.3%
0.82
$\uparrow$6.5%
0.83
$\uparrow$5.1%
Daphnet
0.14s
$\uparrow$42%
0.134s
0.55
$\uparrow$161%
0.44
$\uparrow$42%
0.94
$\uparrow$108%
0.87
$\uparrow$71%
MotionSense
0.08
$\uparrow$142%
0.022s
1
$\uparrow$3.1%
0.99
$\uparrow$2.5%
1
$\uparrow$1.2%
0.99
$\uparrow$1.22%
We expanded our analysis to Daphnet (daphnet_freezing_of_gait_245) and two activities from MotionSense (Malekzadeh:2019:MSD:3302505.3310068) (as a binary classification), in addition to the primary tdcs dataset. The results show an increase in training time across the datasets due to class imbalance handling, while the inference time remained stable. DHWT significantly improved precision, F1-score, and specificity in datasets with class imbalance, confirming its effectiveness across diverse conditions. Daphnet showed a 161% boost in precision and a 108% improvement in specificity. These results support the generalizability of DHWT without compromising computational efficiency (Table 7).
4.1.3. Robustness Across Diverse Subjects
@@
4.2. Clinical Correlation of Predicted FoG Scores
To assess the clinical relevance of LIFT-PD’s predictions, we computed correlation coefficients between the model’s FoG scores and standard clinical metrics. Here, FoG scores were defined as the per-subject proportion of detected FoG windows during test sessions. This metric provides a summary-level indicator of how frequently the model detected FoG-related patterns in each participant. We found that the FoG scores were significantly correlated with NFOGQ scores (Pearson’s r = 0.72, p << 0.01) and UPDRS-III Off scores (r = 0.63, p << 0.05), both widely used clinical markers of disease severity and motor impairment. These results indicate strong alignment between the model’s predictions and clinical assessments, supporting the model’s translational potential for patient monitoring.
4.3. Statistical Analysis of Performance Differences
Table 8. The paired t-test results comparing LIFT-PD to the supervised models
Metric t-statistic p-value Interpretation
Precision 3.18 0.011 Statistically significant improvement with LIFT-PD (p<<0.05)
Recall 1.87 0.094 Not statistically significant (p>>0.05), though LIFT-PD shows a positive trend
F1 Score 4.20 0.0023 Highly significant improvement in F1 score with LIFT-PD.
Specificity 2.74 0.023 Significant improvement in reducing false positives (p<<0.05).
Accuracy 4.27 0.0021 Highly significant improvement in overall accuracy
To statistically validate the improvements observed with LIFT-PD over the supervised baseline, we conducted paired t-tests across multiple performance metrics. The results show statistically significant improvements in precision (p = 0.011), F1-score (p = 0.0023), specificity (p = 0.023), and accuracy (p = 0.0021), confirming that LIFT-PD outperforms the supervised model in key areas relevant to clinical application. Although the improvement in recall (p = 0.094) was not statistically significant at the 5% level, it still showed a positive trend, suggesting consistent performance across diverse patient groups. These findings support the robustness and reliability of the proposed approach.
4.4. Compare with State-of-the-Art Models
Table 9 compares the performance of our LIFT-PD model with state-of-the-art methods for detecting FoG in PD patients. For fairness, all models were implemented from scratch and evaluated on the same dataset using consistent experimental protocols. Although the Multi-head CNN (10.1016/j.artmed.2022.102459) achieves the highest detection rates for FoG episodes (97.27%) and windows (94.64%), its precision (0.545) and specificity (0.491) are low, indicating a high false positive rate. High false positives can reduce the effectiveness of cueing due to patients’ adaptiveness to “always on” interventions (ginis2018cueing) (cubo2004short). The one-class classifier (naghavi2021towards) shows high precision (0.856) and specificity (0.891) but lower recall (0.716), missing many true FoG events, which can lead to inadequate monitoring and delayed interventions. The semi-supervised model (mikos2017real) shows a recall around 12% lower and specificity approximately 13% lower than LIFT-PD, making it less suitable for in-home monitoring with a single accelerometer.
Table 9. Comparison with state-of-the-art models